# Doctor-R1: Mastering Clinical Inquiry with Experiential Agentic Reinforcement Learning

### v1

@@ Abstract
The professionalism of a human doctor in outpatient service depends on two core abilities: the ability to make accurate medical decisions and the medical consultation skill to conduct strategic, empathetic patient inquiry. Existing Large Language Models (LLMs) have achieved remarkable accuracy on medical decision-making benchmarks. However, they often lack the ability to conduct the strategic and empathetic consultation, which is essential for real-world clinical scenarios. To address this gap, we propose Doctor-R1, an AI doctor agent trained to master both of the capabilities by ask high-yield questions and conduct strategic multi-turn inquiry to guide decision-making. Our framework introduces three key components: a multi-agent interactive environment, a two-tiered reward architecture that separately optimizes clinical decision-making and communicative inquiry skills, and an experience repository to ground policy learning in high-quality prior trajectories. We evaluate Doctor-R1 on OpenAI’s HealthBench and MAQuE, assessed across multi-facet metrics, such as communication quality, user experience, and task accuracy. Remarkably, Doctor-R1 surpasses state-of-the-art open-source specialized LLMs by a substantial margin with higher parameter efficiency and outperforms powerful proprietary models. Furthermore, the human evaluations show a strong preference for Doctor-R1 to generate human-preferred clinical dialogue, demonstrating the effectiveness of the framework. For more details refer to https://github.com/thu-unicorn/Doctor-R1.
Figure 1: The Gap Between Static Medical Desion Task and Dynamic Clinical Inquiry: Top-tiers models demonstrate high performance on static medical decision benchmarks. However, these models adhere to a generic script in a high-risk clinical case, failing to adapt to the complex scenario. Doctor-R1 demonstrates a strategic inquiry process, showing the effectiveness of our framework.
@@ 1 Introduction
To assess this capability, a doctor agent should adhere to three key principles: 1) Strategic and Dynamic Inquiry: An effective doctor agent must strategically ask high-yield questions to quickly identify critical risks, rather than following a generic script. For example, in Figure 1, models such as GPT-4.1, HuatuoGPT-o1-70B (chen2024huatuogpto1medicalcomplexreasoning), and DoctorAgent-RL (feng2025doctoragentrlmultiagentcollaborativereinforcement), fails this principle by rigidly following a generic questioning script, rather than adapting its inquiry based on the patient’s responses, resulting in low efficiency and an unsafe consultation. 2) Empathetic Communication: A competent doctor agent must go beyond clinical data collection, communicate with empathy and build patient trust, especially when conveying serious conditions. For instance, Baichuan-M2 (baichuan2025m2) fails this principle when handling emergencies in Table LABEL:tab:inquiry_example. 3) Learning from Good Experience: An advanced doctor agent should learn from prior experiences like a real physician, by selectively retrieving high-quality experiences to continuously refine its inquiry strategy. Existing models fail to support dynamic inquiry (Principle 1), empathic communication (Principle 2), and high-quality experience learning (Principle 3), which relying primarily on similarity-based retrieval (yan2025memoryr1enhancinglargelanguage; zhong2023memorybankenhancinglargelanguage; li2025agenthospitalsimulacrumhospital).
@@
  • Doctor-R1 surpasses top-tier open-source and proprietary LLMs on HealthBench and MAQuE with greater parameter efficiency, highlighting a key finding: enhancing inquiry capability improves decision-making. These results are further validated by strong human preferences in evaluation studies.
@@ 3.1 Dynamic Interactive Environment
This closed-loop feedback system provides the foundation for training the polify of the doctor agent. Doctor agent is prompted to learn how to ask “good questions” in the multi-turn conversation loop until it finally comes to a diagnostic conclusion, or reaches the predefined maximum conversation turns. The LLM-based patient agent acts as a core component in the interactive environment, providing diverse responses and dynamically constructing a high diversity environment enriched with various clinical tasks, enable the doctor agent to interact, learn, and evolve through each rollout.
@@ 4 Evaluation Results
Our primary evaluations are conducted on HealthBench (arora2025healthbenchevaluatinglargelanguage) and MAQuE (MAQuE), two benchmarks featuring complex clincal scenarios and multi-turn dialogues. To validate that Doctor-R1’s specialized training does not compromise its foundational knowledge, we also evaluate it on the static QA benchmark MedQA (jin2020diseasedoespatienthave) and MMLU (hendrycks2021measuringmassivemultitasklanguage), to confirm the robustness of our model’s knowledge base.
@@ 4.1 Overall Performance
Table 1: Overall model performance on HealthBench Main (best results of open-source models are bolded)

### v2

@@ Abstract
The professionalism of a human doctor in outpatient service depends on two core abilities: the ability to make accurate medical decisions and the medical consultation skill to conduct strategic, empathetic patient inquiry. Existing Large Language Models (LLMs) have achieved remarkable accuracy on medical decision-making benchmarks. However, they often lack the ability to conduct the strategic and empathetic consultation, which is essential for real-world clinical scenarios. To address this gap, we propose Doctor-R1, an AI doctor agent trained to master both of the capabilities by ask high-yield questions and conduct strategic multi-turn inquiry to guide decision-making. Our framework introduces three key components: a multi-agent interactive environment, a two-tiered reward architecture that separately optimizes clinical decision-making and communicative inquiry skills, and an experience repository to ground policy learning in high-quality prior trajectories. We evaluate Doctor-R1 on HealthBench and MAQuE, assessed across multi-facet metrics, such as communication quality, user experience, and task accuracy. Remarkably, Doctor-R1 surpasses state-of-the-art specialized LLMs by a substantial margin with higher parameter efficiency and outperforms powerful proprietary models. Human expert evaluations show that Doctor-R1 achieves superior clinical capability and patient-centric performance. For more details refer to https://github.com/thu-unicorn/Doctor-R1.
Figure 1: The Gap Between Static Medical Decision Task and Dynamic Clinical Inquiry: Top-tier models achieve strong results on static benchmarks but show limited adaptability to the complex scenario. Doctor-R1 demonstrates a strategic inquiry process and differential diagnosis, showing the effectiveness of our framework (see Table LABEL:tab:inquiry_example for complete case study).
@@ 1 Introduction
To assess this capability, a doctor agent should adhere to three key principles: 1) Strategic and Dynamic Inquiry: An effective doctor agent must strategically ask high-yield questions and perform differential diagnosis to quickly identify critical risks. For example, in Figure 1, models such as GPT-4.1 and HuatuoGPT-o1-70B (chen2024huatuogpto1medicalcomplexreasoning) may fail this principle by asking non-targeted questions, rather than adapting its inquiry based on the patient’s responses, resulting in low efficiency and an unsafe consultation. 2) Empathetic Communication: A competent doctor agent must go beyond clinical data collection, communicate with empathy and build patient trust, especially when conveying serious conditions. For instance, Baichuan-M2 (baichuan2025m2) and DoctorAgent-RL (feng2025doctoragentrlmultiagentcollaborativereinforcement) fail this principle when handling emergencies in Table LABEL:tab:inquiry_example. 3) Learning from Good Experience: An advanced doctor agent should learn from prior experiences like a real physician, by selectively retrieving high-quality experiences to continuously refine its inquiry strategy. Existing models are limited in supporting dynamic inquiry (Principle 1), empathic communication (Principle 2), and high-quality experience learning (Principle 3), which relying primarily on similarity-based retrieval (yan2025memoryr1enhancinglargelanguage; zhong2023memorybankenhancinglargelanguage; li2025agenthospitalsimulacrumhospital).
@@
  • Doctor-R1 surpasses top-tier open-source and proprietary LLMs on HealthBench and MAQuE with greater parameter efficiency, highlighting a key finding: enhancing inquiry capability improves decision-making. These results are further validated by human expert evaluations, which highlight the superior clinical capability and patient-centric performance of our framework.
@@ 3.1 Dynamic Interactive Environment
This closed-loop feedback system provides the foundation for training the policy of the doctor agent. Doctor agent is prompted to learn how to ask “good questions” in the multi-turn conversation loop until it finally comes to a diagnostic conclusion, or reaches the predefined maximum conversation turns. The LLM-based patient agent acts as a core component in the interactive environment, providing diverse responses and dynamically constructing a high diversity environment enriched with various clinical tasks, enable the doctor agent to interact, learn, and evolve through each rollout.
@@ 4 Evaluation Results
Our primary evaluations are conducted on HealthBench (arora2025healthbenchevaluatinglargelanguage) and MAQuE (MAQuE), two benchmarks featuring complex clincal scenarios and multi-turn dialogues. To validate that our training does not compromise its foundational knowledge, we also evaluate on the static QA benchmark MedQA (jin2020diseasedoespatienthave) and MMLU (hendrycks2021measuringmassivemultitasklanguage).
@@ 4.1 Overall Performance
Table 1: Overall performance on HealthBench Main (best results of open-source models are bolded)